# ViSAGe: A Global-Scale Analysis of Visual Stereotypes in Text-to-Image Generation

### v2

@@ 1 Introduction
  • We publicly release the dataset ViSAGe : Visual Stereotypes Around the Globe where we make a critical distinction between ‘visual’ and ‘non-visual’ stereotypes in images. We identify a list of 385 visual attributes, and also introduce a broad-coverage image dataset with annotations of visual markers of stereotypes present in 40,057 image-attribute pairs, representing different identity groups.  [^1] For access to the dataset, please fill out the form at : https://forms.gle/3kqsry11D4gLKHQc7
@@ 2 Related Work
Cho et al. (2023) show that T2I models reflect specific gender/skin tone biases from training data. Zhang et al. (2023) study gender presentation in T2I models. Bianchi et al. (2023) highlight the presence of stereotypes by using prompts containing attributes and associating the generated visual features with demographic groups. Stable Bias Luccioni et al. (2023) prompts T2I models with a combination of ethnicity/gender and profession and evaluates profession-based stereotypes. Basu et al. (2023) use location-based prompts to quantify geographical representativeness of the generated images. Qadri et al. (2023) identify harmful stereotypes and an ‘outsiders gaze’ in image generation in the South Asian context. We prompt T2I models with ‘identity groups’ around the globe to study the prevalent regional stereotypes in their default visual representation. We identify ‘visual’ stereotypes, related to the concept of ‘imageable’ synsets Yang et al. (2020), to conduct this analysis at scale.
@@ 3.1 Identifying Visual Stereotypes
The annotators are presented with an attribute and a statement of the form of ‘The attribute [attr] can be visually depicted in an image.’ The attribute presented in the sentence is selected from a list of all unique attributes (noun/noun phrases, and adjective/adjective phrases) from SeeGULL. The annotators are required to assign a Likert scale rating to each attribute indicating the extent to which they agree or disagree with the above statement w.r.t. to the given attribute term. We note that the visual nature of attributes lie on a spectrum. While there may be clearly defined attributes on either extremes, the ‘visual’ and ‘non-visual’ attributes in the middle are more subjective. For our subsequent analysis, we exclude all attributes where any annotator expressed uncertainty or disagreement regarding the visual nature. Consequently, we deem terms where all annotators at least agreed about their visual nature, as ‘visual’ resulting in a selection of 385 out of the original 1994 attributes. Please refer to the Appendix 7.1.1 for details.
We recruited annotators based out of the United States, and proficient in English reading and writing. For each attribute, we get annotations from 3 annotators that identify with different geographical origin identities - Asian, European, and North American. Annotators were professional data labelers and were compensated at rates above the prevalent market rates.
@@ 3.2 Detecting Visual Stereotypes in Text-to-Image Generation
Annotating stereotypes in images at scale can often times be resource-intensive and time-consuming. Therefore, we also investigate the feasibility of using automated techniques for stereotype detection in images. Using an already existing image captioning model, CLIP (Contrastive Language-Image Pre-training), we get top 50 captions for each image for the same set of the 15 images per identity group previously generated. Building on the already identified visual stereotypes, we identify the stereotypes in captions by performing a string match. To further understand how uniquely a stereotypical attribute $attr_{s}$ is present in the caption of the images of an identity group, we compute a salience score of the attributes w.r.t. the identity group $S(attr_{s},id)$. We use a modified tf-idf metric as follows: $S(attr_{s},id)=tf(attr_{s},id)\cdot idf(attr_{s},C)$. The function $tf(attr_{s},id)$ represents the smoothed relative frequency of attribute $attr_{s}$ for the identity group $id$; and the function $idf(attr_{s},C)$ denotes the inverse document frequency of the attribute term $attr_{s}$ in all the captions ‘C’, reflecting the importance of the attribute across all the captions. A higher salience score of an attribute, indicates a more unique association of the attribute with the identity group. For each identity group, we then extract the most salient associated visual stereotypes as present in the captions.
@@ 6 Conclusion
Although we cover a wide range of visual stereotypes, annotation about the visual nature of attributes is potentially subjective. While we attempt to account for this subjectivity by (1) using a Likert scale as opposed to binary labels, and (2) capture diversity in our annotator pool w.r.t. gender and geographical region, we recognize that our annotations may still fail to capture other dimensions of subjectivity. Moreover, our evaluation of prevalence of stereotypes in the generated images is limited by the stereotypes present in textual resources like SeeGULL. More regional stereotypes from different textual resources can be included to further expand the overall coverage. We evaluate the images generated using Stable Diffusion V1-4 and believe the results would hold even on other Text-to-Image models, but that needs to be verified in the future work where our analysis may be applied to other generative image models. We chose Stable Diffusion due to its ease of access for academic research, and since our core objective is to demonstrate gaps in existing stereotype bias evaluation approaches, rather than demonstrating biases in any particular or all image generation models. Our work focuses only on regional stereotypes and their presence in the generated images. The proposed framework could be used to experiment with other axes like gender, race, ethnicity, etc., as well. We believe that our approach is extensible to other types of stereotypes, provided such data repositories exist. Future iterations of such data collection and evaluation should take more participatory approach and involve communities with lived experiences on the harms of bias in society.
@@
We evaluate the stereotypes generated in visual depictions of people, based on their identity associated with a geographical location. We acknowledge that this is a complex notion of identity and will have overlap with other facets of a person’s identity, such as race, which is not a focus of this work. We also emphasize that this work does not cover all possible visual depictions of stereotypes and thus, it’s methods or dataset should not be used as a standard to deem model generations free of stereotypes. This work is an initial step towards a systematic, global-scale evaluations of T2I models for stereotypical depictions, and we hope for future work to build on this to attain more coverage of potential stereotypical and harmful depictions in generations.
Acknowledgements
We thank Utsav Prabhu, and Kathy Meier-Hellstern for their feedback and advice.

### v3

@@ 1 Introduction
  • We publicly release the dataset ViSAGe111 https://github.com/google-research-datasets/visage : Visual Stereotypes Around the Globe where we make a critical distinction between ‘visual’ and ‘non-visual’ stereotypes in images. We identify a list of 385 visual attributes, and also introduce a broad-coverage image dataset with annotations of visual markers of stereotypes present in 40,057 image-attribute pairs, representing different identity groups.
@@ 2 Related Work
Cho et al. (2023) show that T2I models reflect specific gender/skin tone biases from training data. Fraser and Kiritchenko (2024) examine gender and racial bias in vision language models. Zhang et al. (2023) study gender presentation in T2I models. Ungless et al. (2023) demonstrate that images of non-cisgender identities are more stereotyped and more sexualised. Bianchi et al. (2023) highlight the presence of stereotypes by using prompts containing attributes and associating the generated visual features with demographic groups. Stable Bias Luccioni et al. (2023) prompts T2I models with a combination of ethnicity/gender and profession and evaluates profession-based stereotypes. Basu et al. (2023) use location-based prompts to quantify geographical representativeness of the generated images. Qadri et al. (2023) identify harmful stereotypes and an ‘outsiders gaze’ in image generation in the South Asian context. We prompt T2I models with ‘identity groups’ around the globe to study the prevalent regional stereotypes in their default visual representation. We identify ‘visual’ stereotypes, related to the concept of ‘imageable’ synsets Yang et al. (2020), to conduct this analysis at scale.
@@ 3.1 Identifying Visual Stereotypes
The annotators are presented with an attribute and a statement of the form of ‘The attribute [attr] can be visually depicted in an image.’ The attribute presented in the sentence is selected from a list of all unique attributes (noun/noun phrases, and adjective/adjective phrases) from SeeGULL. The annotators are required to assign a Likert scale rating to each attribute indicating the extent to which they agree or disagree with the above statement w.r.t. to the given attribute term. We note that the visual nature of attributes lie on a spectrum. While there may be clearly defined attributes on either extremes, the ‘visual’ and ‘non-visual’ attributes in the middle are more subjective. For our subsequent analysis, we exclude all attributes where any annotator expressed uncertainty or disagreement regarding the visual nature. Consequently, we deem terms where all annotators at least agreed about their visual nature, as ‘visual’ resulting in a selection of 385 out of the original 1994 attributes. Please refer to the Appendix 7.1.1 for details. We recruited annotators based out of the United States, and proficient in English reading and writing. For each attribute, we get annotations from 3 annotators that identify with different geographical origin identities - Asian, European, and North American. Annotators were professional data labelers and were compensated at rates above the prevalent market rates.
@@ 3.2 Detecting Visual Stereotypes in Text-to-Image Generation
Annotating stereotypes in images at scale can often times be resource-intensive and time-consuming. Therefore, we also investigate the feasibility of using automated techniques for stereotype detection in images. We use the already existing CLIP (Contrastive Language-Image Pre-training) embeddings combined with the BART model to generate image captions. An image is mapped into the CLIP embedding space, and the embeddings are then scored against a cache of n-grams. The top-k embeddings are then sent to BART to generate candidate captions. We get top 50 captions for each image for the same set of the 15 images per identity group previously generated. Building on the already identified visual stereotypes, we identify the stereotypes in captions by performing a string match. To further understand how uniquely a stereotypical attribute $attr_{s}$ is present in the caption of the images of an identity group, we compute a salience score of the attributes w.r.t. the identity group $S(attr_{s},id)$. We use a modified tf-idf metric as follows: $S(attr_{s},id)=tf(attr_{s},id)\cdot idf(attr_{s},C)$. The function $tf(attr_{s},id)$ represents the smoothed relative frequency of attribute $attr_{s}$ for the identity group $id$; and the function $idf(attr_{s},C)$ denotes the inverse document frequency of the attribute term $attr_{s}$ in all the captions ‘C’, reflecting the importance of the attribute across all the captions. A higher salience score of an attribute, indicates a more unique association of the attribute with the identity group. For each identity group, we then extract the most salient associated visual stereotypes as present in the captions.
@@ 6 Conclusion
Acknowledgements
We thank Kathy Meier-Hellstern, and Utsav Prabhu for their detailed feedback and Susan Hao for useful discussions. We also thank the anonymous reviewers for the constructive feedback during the review process.
@@ Acknowledgements
Intended Usage
Not all attributes are undesirable in a T2I generation context. Determining whether a T2I model should reflect certain attributes associated with specific identity groups is a non-trivial question and can be most accurately determined by the respective developer and the downstream use case, taking into account the type of harm Shelby et al. (2023) experienced by the end user. Our objective in this work is not to advocate for a stereotype-neutral generation but to equip practitioners with resources and evaluation methods to assesss, on a scalable and effective basis, the extent to which global stereotypes are reproduced in model generations. We also aim to facilitate an understanding of the nature of stereotypes, including their potential offensiveness, through this work. We hope this enables model builders and platforms to better prioritize and appropriately intervene in their model generations based on their use case.
Limitations
Although we cover a wide range of visual stereotypes, annotation about the visual nature of attributes is potentially subjective. While we attempt to account for this subjectivity by (1) using a Likert scale as opposed to binary labels, and (2) capture diversity in our annotator pool w.r.t. gender and geographical region, we recognize that our annotations may still fail to capture other dimensions of subjectivity. Moreover, our evaluation of prevalence of stereotypes in the generated images is limited by the stereotypes present in textual resources like SeeGULL. More regional stereotypes from different textual resources can be included to further expand the overall coverage. We evaluate the images generated using Stable Diffusion V1-4 and believe the results would hold even on other Text-to-Image models, but that needs to be verified in the future work where our analysis may be applied to other generative image models. We chose Stable Diffusion due to its ease of access for academic research, and since our core objective is to demonstrate gaps in existing stereotype bias evaluation approaches, rather than demonstrate biases in any particular or all image generation models. Our work focuses only on regional stereotypes and their presence in the generated images. The proposed framework could be used to experiment with other axes like gender, race, ethnicity, etc., as well. We believe that our approach is extensible to other types of stereotypes, provided such data repositories exist. Future iterations of such data collection and evaluation should take more participatory approach and involve communities with lived experiences on the harms of bias in society.
@@
We evaluate stereotypes in the generated visual depictions of people, based on their identity associated with a geographical location. We acknowledge that this is a complex notion of identity and overlaps with other facets of a person’s identity, such as race, which is not a focus of this work. We also emphasize that this work does not cover all possible visual depictions of stereotypes and thus, it’s methods or dataset should not be used as a standard to deem model generations free of stereotypes. This is an initial step towards a systematic, global-scale evaluations of T2I models for stereotypical depictions, and we hope for future work to build on this to attain more coverage of potential stereotypical and harmful depictions in generations.